Chapter 1

Too much vspace

The user interacts with a proof assistant via a proof script in the following general sense.

Definition 1.0.1 (proof script) A proof script is the sequence of instructions a user gives a proof assistant to prove a statement. The user interacts with a proof assistant via a proof script in the following general sense Definition 1.0.2 (proof script) A proof script is the sequence of instructions a user gives a proof assistant to prove a statement. The user interacts with a proof assistant via a proof script in the following general sense.

Definition 1.0.3 (proof script) A proof script is the sequence of instructions a user gives a proof assistant to prove a statement. The user interacts with a proof assistant via a proof script in the following general sense Definition 1.0.4 (proof script) A proof script is the sequence of instructions a user gives a proof assistant to prove a statement.